OpenStudy (anonymous):

Its about constant acceleration stuff (m1) While acceleration at 1ms^-2 for 4 sec, a particle travels a distance of 28m. What is its initial velocity???

OpenStudy (anonymous):

s = ut + 0.5at^2 28 = 4u + 0.5 x 1 x 16 28=4u + 8 4u=20 u=5

OpenStudy (anonymous):

THANKS!!! :)